Suppose the sides of a cube are each increased by 20%.
What is the percentage increase in the surfaace area of the solid?
What is the percentage increase in the volume of the solid?

Answer by Alan3354(69443)

Suppose the sides of a cube are each increased by 20%.
What is the percentage increase in the surfaace area of the solid?
Area is a function of the square of the sides.
1.2*s = 1.2^2*area
= 1.44*area
= 44% increase
----------------
What is the percentage increase in the volume of the solid?
Volume is a function of the cube of the sides.
1.2*s = 1.2^3*V
= 1.728V
= 72.8% increase